Solve the equations.
3t(t+5)-t2=2t2+4t-1

Answer by umail08(87)

3t(t+5)-t2 = 2t2+4t-1
3t2+15t-t2=2t2+4t-1
2t2+15t=2t2+4t-1
2t2-2t2+15t=2t2-2t2+4t-1
15t=4t-1
15t-4t=4t-4t-1
11t=-1
11t/11 = -1/11
t = -1/11